# Towards Measuring and Modeling “Culture” in LLMs: A Survey

### v2

@@ Abstract
We present a survey of 39 recent papers that aim to study cultural representation and inclusion in large language models. We observe that none of the studies define “culture,” which is a complex, multifaceted concept; instead, they probe the models on some specially designed datasets which represent certain aspects of “culture.” We call these aspects the proxies of cultures, and organize them across three dimensions of demographic, semantic and linguistic-cultural interaction proxies. We also categorize the probing methods employed. Our analysis indicates that only certain aspects of “culture,” such as values and objectives, have been studied, leaving several other interesting and important facets, especially the multitude of semantic domains Thompson et al. (2020) and aboutness Hershcovich et al. (2022), unexplored. Two other crucial gaps are the lack of robustness and situatedness of the current methods. Based on these observations, we provide several recommendations for a holistic and practically useful research agenda for furthering cultural inclusion in LLMs and LLM-based applications.
@@ 1 Introduction
In this work, we survey 39 NLP papers that study cultural representation, awareness or bias in LLMs either explicitly (e.g., Huang and Yang (2023a), Zhou et al. (2023), Cao et al. (2024)) or implicitly (e.g., Wan et al. (2023)). It is quickly apparent that these papers either do not attempt to define culture or where they do, use very high-level definitions. For example, a common definition is “the way of life of a collective group of people, [that] distinguishes them from other groups with other cultures” (Mora, 2013; Shweder et al., 2007; Hershcovich et al., 2022). Not only do the papers typically use broad-brush definitions, most do not engage in a critical discussion on the topic. [^1] The situation is similar to that described in  Blodgett et al. (2020) in the context of research on “bias”. This is perhaps unsurprising as “culture” is a concept which evades simple definition. It is multifaceted, meaning different things to different people at different times. For example, some of the many and often implicitly applied meanings of culture include a.) “Cultural Heritage” such as art, music, and food habits [^2] https://uis.unesco.org/sites/default/files/documents/analysis_sdg_11.4.1_2022_final_alt_cover_0.pdf Blake (2000), b.) “Interpersonal Interactions” between people from different backgrounds (e.g., ways of speaking in a meeting, politeness norms) Monaghan et al. (2012), or c.) The “Ways of Life” of a collective group of people distinguishing them from other groups. There are a variety of sociological descriptions of culture, for example, Parsons (1972) describes it as the the pattern of ideas and principles which abstractly specify how people should behave, but which do so in ways which prove practically effective relative to what people want to do Münch et al. (1992); Parsons (1972). However, these too are high-level and hard to concretise. Further complications arise because the instantiation of culture is necessarily situated. Every individual and group lies at the intersection of multiple cultures (defined by their political, professional, religious, regional, class-based and other affiliations) and these are invoked according to the situation, typically in contrast to another group(s).
@@ 2.1 Identification and Categorization of Relevant Papers
We read these papers and labeled them for (a) the definition of culture subscribed to in the paper, (b) the method used for probing the LLM for cultural awareness/bias, and (c) the languages and the cultures (thus defined) that were studied. It became apparent during the annotation process that none of the papers attempted to explicitly define “culture.” Therefore, instead of a definition, we decided to label the papers for (1) the proxy for culture used in a study and (2) the aspects of language-culture interaction that were studied. Using these labels, we then built taxonomies of the object of study and the method bottoms-up, which are described in this section, and also used to organize the survey findings in Sections 3 and 4.
@@ 2.2.1 Proxies of Culture
2.2.2 Language-Culture Interaction Axes
@@ 3 Findings: Defining Culture
for tree= grow=east, growth parent anchor=west, parent anchor=east, child anchor=west, calign = edge midpoint, [Culture, root, l sep=2mm,s sep=10mm, [Linguistic-Culture Interaction, xnode, l sep=2mm, [Objectives and Values, tnode, l sep = 2mm, [Kovač et al. (2023);Cao et al. (2023);Johnson et al. (2022);Dev et al. (2023);Gupta et al. (2024);Wan et al. (2023);Tanmay et al. (2023);An et al. (2023);Sorensen et al. (2023);Durmus et al. (2023);, wnode]] [Common Ground, tnode, l sep = 2mm, [Koto et al. (2023);Wibowo et al. (2023);Wang et al. (2023);Johnson et al. (2022);Quan et al. (2020);Zhang et al. (2023);Bauer et al. (2023);Naous et al. (2023);Wu et al. (2023);Nguyen et al. (2023);Huang and Yang (2023a);,wnode] ] [Aboutness, tnode, l sep = 2mm, [, wnode]]] [Demographic Proxies, xnode, l sep=2mm, [Region, tnode, l sep=2mm, [Hershcovich et al. (2022);Koto et al. (2023);Wibowo et al. (2023);Wang et al. (2023);Johnson et al. (2022);Wan et al. (2023);An et al. (2023);Zhang et al. (2023);Durmus et al. (2023);Jha et al. (2023);Ramezani and Xu (2023);Kabra et al. (2023);Zhou et al. (2023);Mukherjee et al. (2023);CH-Wang et al. (2023);Dev et al. (2023);Khanuja et al. (2023);Santy et al. (2023);Cao et al. (2023);Dwivedi et al. (2023),wnode]] [Language, tnode,l sep=2mm, [Hershcovich et al. (2022);Koto et al. (2023);Kovač et al. (2023);Wibowo et al. (2023);Cao et al. (2023);Johnson et al. (2022);Tanmay et al. (2023);Huang and Yang (2023a);Zhang et al. (2023);Kabra et al. (2023);Naous et al. (2023);Shaikh et al. (2023);Zhou et al. (2023);Mukherjee et al. (2023);CH-Wang et al. (2023);Dev et al. (2023);Khanuja et al. (2023);Santy et al. (2023);Das et al. (2023);,wnode]] [Gender, tnode,l sep=2mm, [Cao et al. (2023);Johnson et al. (2022);Wan et al. (2023);An et al. (2023);Wu et al. (2023);,wnode]] [Race,tnode,l sep=2mm, [Johnson et al. (2022);Durmus et al. (2023);,wnode]] [Religion, tnode,l sep=2mm, [Koto et al. (2023);Durmus et al. (2023);Bauer et al. (2023);Das et al. (2023);,wnode]] [Education, tnode,l sep=2mm, [Koto et al. (2023);Quan et al. (2020);Bauer et al. (2023);Shaikh et al. (2023);Wu et al. (2023);Santy et al. (2023);,wnode]] [Ethnicity, tnode,l sep=2mm, [Koto et al. (2023);Johnson et al. (2022);Wan et al. (2023);Durmus et al. (2023);Santy et al. (2023);Cao et al. (2023);Das et al. (2023);,wnode] ]] [Semantic Proxies, xnode,l sep=2mm, [Names,tnode,l sep=2mm, [Wu et al. (2023);Quan et al. (2020);,wnode]] [Basic Actions and Technology,tnode,l sep=2mm, [Durmus et al. (2023);Bauer et al. (2023);,wnode]] [Social and Political Relations,tnode,l sep=2mm, [Wang et al. (2023);Johnson et al. (2022);Durmus et al. (2023);Shaikh et al. (2023);Bauer et al. (2023);Feng et al. (2023);,wnode]] [Food and Drink,tnode,l sep=2mm, [Palta and Rudinger (2023);Cao et al. (2024);,wnode]] [Emotions and Values,tnode,l sep=2mm, [Hershcovich et al. (2022);Kovač et al. (2023);Koto et al. (2023);Wibowo et al. (2023);Cao et al. (2023);Johnson et al. (2022);Wan et al. (2023);Tanmay et al. (2023);Zhang et al. (2023);Shaikh et al. (2023);Jiang et al. (2022);Talat et al. (2021);Huang and Yang (2023b);Naous et al. (2023);Wu et al. (2023);Fung et al. (2023);Mukherjee et al. (2023);Santy et al. (2023);Cao et al. (2024);,wnode]] ]]
@@ 3.2 Semantic Proxies
Of the other semantic domains, Food and Beverages has been studied by Palta and Rudinger (2023) where a manually-curated set of CommonsenseQA-style questions focused on food-related customs is developed for probing cultural biases in commonsense reasoning systems; and by Cao et al. (2024), who introduce CulturalRecipes – a cross-cultural recipe adaptation dataset in Mandarin Chinese and English, highlighting culinary cultural exchanges.
@@ 3.3 Linguistic-Culture Interaction
Huang and Yang (2023a); Wibowo et al. (2023); Koto et al. (2023); Zhang et al. (2023) evaluate language models’ understanding of common ground through natural language inference (NLI) and commonsense reasoning tests, as they assess the models’ awareness of shared cultural contexts. SeaEval Wang et al. (2023) proposes 4 datasets to probe the culture-specific knowledge in LLMs. EtiCor Dwivedi et al. (2023) is a large dataset of social norms or etiquette collected from do’s and don’ts sections of travel websites. Naous et al. (2023) prompts LLMs in Arabic and English to test for common grounds in Arabic or Islamic cultures. Johnson et al. (2022) explores common ground by examining socio-political policies and norms embedded in non-American cultures.

### v3

@@ Abstract
We present a survey of 55 recent papers that aim to study cultural representation and inclusion in large language models. We observe that none of the studies define “culture,” which is a complex, multifaceted concept; instead, they probe the models on some specially designed datasets which represent certain aspects of “culture.” We call these aspects the proxies of cultures, and organize them across three dimensions of demographic, semantic and linguistic-cultural interaction proxies. We also categorize the probing methods employed. Our analysis indicates that only certain aspects of “culture,” such as values and objectives, have been studied, leaving several other interesting and important facets, especially the multitude of semantic domains Thompson et al. (2020) and aboutness Hershcovich et al. (2022), unexplored. Two other crucial gaps are the lack of robustness and situatedness of the current methods. Based on these observations, we provide several recommendations for a holistic and practically useful research agenda for furthering cultural inclusion in LLMs and LLM-based applications. Compilation and details of papers used for the survey can be found via our GitHub repository [^1] https://github.com/faridlazuarda/cultural-llm-papers
@@ 1 Introduction
In this work, we survey 39 NLP papers that study cultural representation, awareness or bias in LLMs either explicitly (e.g., Huang and Yang (2023a), Zhou et al. (2023), Cao et al. (2024b)) or implicitly (e.g., Wan et al. (2023)). It is quickly apparent that these papers either do not attempt to define culture or where they do, use very high-level definitions. For example, a common definition is “the way of life of a collective group of people, [that] distinguishes them from other groups with other cultures” (Mora, 2013; Shweder et al., 2007; Hershcovich et al., 2022). Not only do the papers typically use broad-brush definitions, most do not engage in a critical discussion on the topic. [^2] The situation is similar to that described in  Blodgett et al. (2020) in the context of research on “bias”. This is perhaps unsurprising as “culture” is a concept which evades simple definition. It is multifaceted, meaning different things to different people at different times. For example, some of the many and often implicitly applied meanings of culture include a.) “Cultural Heritage” such as art, music, and food habits [^3] https://uis.unesco.org/sites/default/files/documents/analysis_sdg_11.4.1_2022_final_alt_cover_0.pdf Blake (2000), b.) “Interpersonal Interactions” between people from different backgrounds (e.g., ways of speaking in a meeting, politeness norms) Monaghan et al. (2012), or c.) The “Ways of Life” of a collective group of people distinguishing them from other groups. There are a variety of sociological descriptions of culture, for example, Parsons (1972) describes it as the the pattern of ideas and principles which abstractly specify how people should behave, but which do so in ways which prove practically effective relative to what people want to do Münch et al. (1992); Parsons (1972). However, these too are high-level and hard to concretise. Further complications arise because the instantiation of culture is necessarily situated. Every individual and group lies at the intersection of multiple cultures (defined by their political, professional, religious, regional, class-based and other affiliations) and these are invoked according to the situation, typically in contrast to another group(s).
@@ 2.1 Identification and Categorization of Relevant Papers
We read these papers and labeled them for (a) the definition of culture subscribed to in the paper, (b) the method used for probing the LLM for cultural awareness/bias, and (c) the languages and the cultures (thus defined) that were studied. It became apparent during the annotation process that none of the papers attempted to explicitly define “culture.” Therefore, instead of a definition, we decided to label the papers for (1) the proxy for culture used in a study and (2) the aspects of linguistic-culture Interaction that were studied. Using these labels, we then built taxonomies of the object of study and the method bottoms-up, which are described in this section, and also used to organize the survey findings in Sections 3 and 4.
@@ 2.2.1 Proxies of Culture
2.2.2 Linguistic-Culture Interaction Axes
@@ 3 Findings: Defining Culture
for tree= grow=east, growth parent anchor=west, parent anchor=east, child anchor=west, calign = edge midpoint, [Linguistic-Culture Interaction, xnode, l sep=2mm, [Objectives and Values, tnode, l sep = 2mm, [Kovač et al. (2023);Cao et al. (2023);Johnson et al. (2022);Dev et al. (2023);Gupta et al. (2024);Wan et al. (2023);Tanmay et al. (2023);An et al. (2023);Sorensen et al. (2023);Durmus et al. (2023);Cao et al. (2024a);, wnode]] [Common Ground, tnode, l sep = 2mm, [Koto et al. (2023);Wibowo et al. (2023);Wang et al. (2023);Johnson et al. (2022);Quan et al. (2020);Zhang et al. (2023);Bauer et al. (2023);Naous et al. (2023);Wu et al. (2023);Nguyen et al. (2023);Huang and Yang (2023a);Liu et al. (2024);,wnode] ] [Aboutness, tnode, l sep = 2mm, [Koto et al. (2024), wnode]]]
{forest}
for tree= grow=east, growth parent anchor=west, parent anchor=east, child anchor=west, calign = edge midpoint, [Demographic Proxies, xnode, l sep=2mm, [Region, tnode, l sep=2mm, [Hershcovich et al. (2022);Koto et al. (2023);Wibowo et al. (2023);Wang et al. (2023);Johnson et al. (2022);Wan et al. (2023);An et al. (2023);Zhang et al. (2023);Durmus et al. (2023);Jha et al. (2023);Ramezani and Xu (2023);Kabra et al. (2023);Zhou et al. (2023);Mukherjee et al. (2023);CH-Wang et al. (2023);Dev et al. (2023);Khanuja et al. (2023);Santy et al. (2023);Cao et al. (2023);Dwivedi et al. (2023);Koto et al. (2024);Cao et al. (2024a);Liu et al. (2024);,wnode]] [Language, tnode,l sep=2mm, [Hershcovich et al. (2022);Koto et al. (2023);Kovač et al. (2023);Wibowo et al. (2023);Cao et al. (2023);Johnson et al. (2022);Tanmay et al. (2023);Huang and Yang (2023a);Zhang et al. (2023);Kabra et al. (2023);Naous et al. (2023);Shaikh et al. (2023);Zhou et al. (2023);Mukherjee et al. (2023);CH-Wang et al. (2023);Dev et al. (2023);Khanuja et al. (2023);Santy et al. (2023);Das et al. (2023);Cao et al. (2024a);,wnode]] [Gender, tnode,l sep=2mm, [Cao et al. (2023);Johnson et al. (2022);Wan et al. (2023);An et al. (2023);Wu et al. (2023);,wnode]] [Race,tnode,l sep=2mm, [Johnson et al. (2022);Durmus et al. (2023);,wnode]] [Religion, tnode,l sep=2mm, [Koto et al. (2023);Durmus et al. (2023);Bauer et al. (2023);Das et al. (2023);,wnode]] [Education, tnode,l sep=2mm, [Koto et al. (2023);Quan et al. (2020);Bauer et al. (2023);Shaikh et al. (2023);Wu et al. (2023);Santy et al. (2023);,wnode]] [Ethnicity, tnode,l sep=2mm, [Koto et al. (2023);Johnson et al. (2022);Wan et al. (2023);Durmus et al. (2023);Santy et al. (2023);Cao et al. (2023);Das et al. (2023);Koto et al. (2024);Liu et al. (2024);,wnode] ]]
{forest}
for tree= grow=east, growth parent anchor=west, parent anchor=east, child anchor=west, calign = edge midpoint, [Semantic Proxies, xnode,l sep=2mm, [Names,tnode,l sep=2mm, [Wu et al. (2023);Quan et al. (2020);,wnode]] [Basic Actions and Technology,tnode,l sep=2mm, [Durmus et al. (2023);Bauer et al. (2023);,wnode]] [Social and Political Relations,tnode,l sep=2mm, [Wang et al. (2023);Johnson et al. (2022);Durmus et al. (2023);Shaikh et al. (2023);Bauer et al. (2023);Feng et al. (2023);Koto et al. (2024);,wnode]] [Food and Drink,tnode,l sep=2mm, [Palta and Rudinger (2023);Cao et al. (2024b);Koto et al. (2024);,wnode]] [Emotions and Values,tnode,l sep=2mm, [Hershcovich et al. (2022);Kovač et al. (2023);Koto et al. (2023);Wibowo et al. (2023);Cao et al. (2023);Johnson et al. (2022);Wan et al. (2023);Tanmay et al. (2023);Zhang et al. (2023);Shaikh et al. (2023);Jiang et al. (2022);Talat et al. (2021);Huang and Yang (2023b);Naous et al. (2023);Wu et al. (2023);Fung et al. (2023);Mukherjee et al. (2023);Santy et al. (2023);Cao et al. (2024b);Cao et al. (2024a);Liu et al. (2024);,wnode]] ]
@@ 3.2 Semantic Proxies
Of the other semantic domains, Food and Beverages has been studied by Palta and Rudinger (2023) where a manually-curated set of CommonsenseQA-style questions focused on food-related customs is developed for probing cultural biases in commonsense reasoning systems; and by Cao et al. (2024b), who introduce CulturalRecipes – a cross-cultural recipe adaptation dataset in Mandarin Chinese and English, highlighting culinary cultural exchanges.
@@ 3.3 Linguistic-Culture Interaction
Huang and Yang (2023a); Wibowo et al. (2023); Koto et al. (2023); Zhang et al. (2023) evaluate language models’ understanding of common ground through natural language inference (NLI) and commonsense reasoning tests, as they assess the models’ awareness of shared cultural contexts. SeaEval Wang et al. (2023) proposes 4 datasets to probe the culture-specific knowledge in LLMs. EtiCor Dwivedi et al. (2023) is a large dataset of social norms or etiquette collected from do’s and don’ts sections of travel websites. Naous et al. (2023) prompts LLMs in Arabic and English to test for common grounds in Arabic or Islamic cultures. Johnson et al. (2022) explores common ground by examining socio-political policies and norms embedded in non-American cultures. Fung et al. (2023) explores discovery of social norms in a multi-cultural and multi-lingual setup from conversations.